Chapter VI

CONCLUDING REMARKS

174

Chapter VI

CONCLUDING REMARKS

Abstract

In this final chapter, I summarize the major conclusions from the previous four chapters concerning the fate of gas during mergers between disk galaxies. This chapter should not be considered a summary of the main results of the thesis, which are given at the end of each chapter. Rather, it is my assessment of the thesis topic. I address each of the questions posed in Chapter I, and briefly present what I believe are the most important avenues of research for the future.

1. The Fate of Gas during Disk-Disk Mergers

There are 21 bullets at the ends of Chapters II–IV which contain the main results of this thesis. Restricting myself to the named topic of the thesis, “The Fate of Gas in Merging Disk Galaxies”, I reduced these to the following four bullets:

• Along the sequence, a larger fraction of the H i appears outside the optical bodies, in extended tidal features. In the final stage all of the H i is found in the tails, with no gas detected within the remnant body. This is due to the depletion of the inner H i rather than its ejection into the tails. The inner H i content decreases in concert with the destruction of the optical disks, suggesting that the process of violent relaxation converts atomic gas into other forms (stars, molecules, hot plasma) as it redistributes the starlight.

• We observed two merger remnants, NGC 3921 and NGC 7252, characterized by r^{1/4} light profiles, atomic gas-poor remnant bodies, a post-burst stellar population, and atomic gas-rich tidal tails. They also have large quantities of central molecular gas and a significant amount of optical fine structure, both of which should decrease over the next few Gyr. These properties support the idea that some spiral galaxies can merge and evolve into ellipticals.

• In both late stage mergers, some of the outer gas is found to be falling back toward the remnant body. This illustrates that much of the tidal material is not in free expansion, but remains bound to the remnants. Dynamical modeling of NGC 7252 shows that initially luminous tails are raised but most of this material falls back before the final merger. The more distant tail material continues to rain back upon the remnant for several more Gyr, in ever decreasing amounts and to ever larger radii. That the remnant remain H i poor suggest that the presently returning gas is continuously converted into other forms. One-fifth of the current tail material is expected to be lost from the remnant.

• The tail kinematics are determined by gravity. Star formation occurs within the tidal features at all stages, making them lively arenas. At the later stages, H ii concentrations coincide with H i concentrations, leading us to believe that self-gravity is important within the tails. If these clumps are bound, they serve to support the idea originally proposed by Zwicky (1956), and verified in recent numerical simulations, that dwarf galaxies can form from tidally ejected material.

2. Answers to the Posed Questions

In section 3 of Chapter I, I presented the eight questions that I wished to answer in this thesis. I now return to each of these questions, drawing from the preceding text for their answers.

(1) Do these systems indeed appear to be spirals or former spirals colliding? If so, does it seem likely that the end product will be an elliptical?

Since the double-tailed morphological criterion used by Toomre (1977) is now taken as the signature of merging disks, the first part of this question really didn’t need to be asked. It was thought that the spiral nature of the progenitors would be further reaffirmed by the presence of atomic gas rich tails, but this was only the case for three of the five systems (Arp 295, NGC 4676, and NGC 7252). In NGC 520 and NGC 3921, at least one atomic gas rich progenitor (i.e. late type spiral) is indicated by a gas-rich tidal feature, but the second tidal appendage is gas poor, which I take as evidence that the second progenitor was an early type spiral (§II.4.2.3-4, §II.6.2.3). The smooth trend in velocities of the gas-rich features is further confirmation of their gravitational origin (§II.6.2), so yes, these five objects do appear to be merging or merged disk galaxies.

The late stage mergers NGC 3921 and NGC 7252 are already classified as S0s, have light profiles typical of ellipticals (Schweizer 1982, Stanford & Bushouse 1991; see also §II.6.3.2) and obey the Faber-Jackson (1976) relationship (Lake & Dressler 1986). NGC 7252 also fits within the fundamental plane (Djorgovski & Davis 1987) on which ellipticals and bulges lie (§III.4.4, Append. VI). The remaining question was “will these systems evolve into ellipticals also in their gas content?”. The low central atomic gas content of both of these systems (§II.5.3), despite the streams of returning atomic gas (§II.6.2.1; Ch. IV), argues for an affirmative answer to this question. However, the evidence is not fully conclusive. It is still possible that these systems will never rid themselves of their central molecular gas and any returning H i, or that this gas will be converted into a noticeable stellar disk, in which case they would remain S0s. In order to more fully constrain the evolutionary fates of these remnants, more objects in the so-called “King Gap” between merger remnants and full fledged ellipticals (§I.2.2.2) need to be studied. This is one of the primary avenues of research I discuss in §3 below.

§2. Answers to the Posed Questions 177

One of the more interesting aspects of this question is how do the mergers rid themselves of their ample atomic gas content? In this thesis I argue that the central H i content decreases in concert with the destruction of the optical disks (§II.5.3). Since it is the process of violent relaxation which redistributes the stellar light into the relaxed luminosity profile typical of ellipticals (Lynden-Bell 1967, van Albada 1982), I conclude that this process is also responsible for an efficient conversion of H i into other phases within the remnant body (§II.6.1.3; §IV.5.1). During violent relaxation gas will move off of its originally non-intersecting orbits, collide/compress and dissipate, and obtaining high densities as it sinks to the center of the potential. I suggest that in the resulting collisions and/or compressions, much of the atomic gas is converted into other forms (molecules, stars, and plasma). Any remaining (or returning) atomic gas may be heated by the resulting central starburst.

An interesting corollary to this conclusion is that if the violent relaxation is “sufficiently” incomplete, the gas may retain much of its original angular momentum and resettle into a galactic-scale disk, which may subsequently form stars. In this case, the merger of two spirals would result in another spiral. Such may be the fate of the NGC 520 system (§II.6.3.4). I find this corollary attractive because it offers a natural solution to a personal apprehension regarding the Merger Hypothesis, which is: if mergers between spirals inevitably lead to ellipticals, we must be living in a “special” epoch in which we are near examples of both types of galaxies. While there are other possible solutions to this objection (e.g. an environmental dependence on merging), I find the aspect of mergers creating the whole range of the Hubble sequence much more satisfying.

(2) Do the chosen systems in fact form sequence? If so, are there any trends along the sequence?

While the five systems chosen for study do form a sequence of mergers arranged according to their time before or since the coalescence of the progenitors main bodies (§II.2), they do not form a sequence in which one system is likely to evolve into another. This is because the mass ratios and Hubble types of the progenitors as well as the encounter geometries are too diverse along the sequence (§II.5). However, we have discovered some “trends” which we believe are important. The first and by far most important was mentioned above: the low central atomic gas content of remnants with relaxed light profiles, despite the evidence for atomic gas rich progenitors.

A second trend is for plumes of ionized gas to appear along the minor axis of the pre-merger systems with the highest gas fraction (§II.5.1). These are believed to be examples of the galactic “blow-out” phenomena (§II.6.1.2). Whether or not blow-out occurs during a merger may have important repercussions for the future evolution of the starburst and/or central gas content. For instance, it has been suggested that the central star bursts are triggered when molecular clouds in pressure equilibrium with the ISM in the outer regions are sent into the higher pressure inner regions by the encounter, where the overpressure causes them to collapse and form stars (Jog & Solomon 1992, Jog & Das 1992). In this scenario, galaxies which have experienced blow-out will vent much of the high pressure inner gas into the halo, and we would expect less violent future star burst than in a system which has retained this high pressure medium.

A final trend is that the optical tails extend beyond the ends of the gaseous tails in the early stage mergers, while the opposite is true in the late stages (§II.5.4). I suggest that the early-stage extensions are due to a halo formed from high velocity dispersion disk stars. These stars will initially make larger radial excursions than the lower-dispersion gas, but either fall back or fade by the late-stages. I look forward to addressing this explanation with detailed numerical models and deep color photometry.

(3) Are there any differences between the gas and stars within these systems? Can we understand them?

The most striking difference between the gas and stars occurs in the late stage systems, NGC 3921 and NGC 7252, where the light is distributed in r^{1/4} profiles while the gas is relegated to the tidal tails. As mentioned in the answer to question #1 above, this is interpreted as evidence that the atomic gas has been converted into other forms within the remnant body during the merger. The symmetry of the tidal forces requires that at least as much disk material is sent inward as is ejected into a tail (§II.6.1.3; §IV.5.1). Since we do not detect any atomic gas in the central regions, several times 10^{9} h^{−2} M_{⊙} of H i must have been converted to other phases during the merger. In both NGC 3921 and NGC 7252 there is evidence for molecular gas, an intermediate aged stellar population, and in NGC 7252 a hot gas phase (§II.6.1.3, §III.4.3). I believe all three are likely sinks, but the data does not allow us to derive a detailed gas budget.

It is reassuring that hydrodynamical simulations yield a similar result (Barnes & Hernquist 1991, 1995). In these simulations, the gas within the main body of the final remnants has three components: a very dense knot of gas containing ∼half of the original disk gas supply concentrated within the central kpc; a diffuse, virialized component (T∼10^{6} K) distributed throughout the remnant body; and dynamically cold streams of gas falling back in from the tails. At the densities of the first phase, atomic gas is expected to be compressed into molecular form; at the temperatures of the second atomic gas would be heated into a X-ray emitting plasma. An important outstanding question is how are the post-merger remnants kept free of the returning atomic gas, despite an accretion rate of 1–2 M_{⊙} yr^{−1} (§IV.5.1).

For the low surface brightness features, there are other fascinating differences between the gas and light. These are discussed in §II.6.2.3. One of the more interesting discoveries is that features with different optical tidal morphologies have different gas characteristics: the edge-brightened tails are gas-rich and the featureless plumes and halos are gas-poor. I believe that these properties are important clues to the identity of the progenitors, arguing that different types of disks will make different types of tidal features. For example, gas-rich disks containing a significant low velocity dispersion stellar component may make tails, while gas poor disks without such a component may make plumes. It is important to test this hypothesis with detailed simulations, as it provides a possible way investigate the pre-merger populations.

We also find displacements between the gas and light in some tidal features which we are at a loss to explain. It therefore remains an important question as to whether there are processes which can separate gas from stars within the tidal features.

Finally, we find that beyond the end of the optical tidal features in NGC 520, NGC 3921, and NGC 7252 there are gaseous extensions which contain significant quantities of atomic hydrogen (M_{HI}≈ 2, 1, 0.4 ×10^{9} h^{−2} M_{⊙}, respectively) at large projected radii (r ≈ 65, 50, 100 h^{−1} kpc, respectively). We believe that these extensions reflect a larger radial distribution of atomic gas in the progenitors (§II.6.2.3). These features may be related to such intriguing purely gaseous structures as the Leo Ring (M_{HI}=1.7×10^{9} M_{⊙}, r=100 kpc; Schneider et al. 1989), the H i plumes in Arp 143 (M_{HI}=2×10^{9} h^{−2} M_{⊙}, r=150 h^{−1} kpc; Appleton et al. 1987) and Arp 215 (M_{HI}=10^{9} h^{−2} M_{⊙}, r=40 h^{−1} kpc; Smith 1991), and the extended H i disk in Mkn 348 (M_{HI}=8×10^{9} h^{−2} M_{⊙}, r=130 h^{−1} kpc; Simkin et al. 1987). These studies underline the unique ability of 21cm mapping observations to investigating such fascinating objects.

(4) What are the substructures within the tails? Can we understand why they are there and/or how they will evolve?

Prominent H i clumps coincident with Hα and stellar enhancements and increased 21cm line widths are seen within the tidal tails (§II.6.2.2, §III.3.3.2, Append. VIII). These clumps resemble dwarf galaxies in their H i contents and blue luminosities and, if bound, in their masses and mass-to-light ratios (§II.6.2.2, §III.4.5). Recent unsuccessful attempts to detect CO in some of these clumps show that they may resemble dwarfs also in their low CO abundance (Smith & Higdon 1994). If such clumps are bound, they serve to support the idea originally proposed by Zwicky (1956), and verified in recent numerical simulations, that dwarf galaxies can form from tidally ejected material. The recent thesis of Duc (1995) studies such clumps in detail. That work provides perhaps the best evidence to date that these clumps are indeed dwarfs-in-formation, by showing that 14 tidal condensations, including the two clumps in the tails of NGC 7252, fall within the same region of the surface brightness–absolute magnitude diagram as dwarf irregulars.

The NGC 7252 simulation predicts that the evolution of the tidal regions will be non-uniform, with self-gravitating regions gaining material at the expense of adjacent areas. Such clumps should therefore eventually become distinct from the surrounding tidal material (§IV.5.2). This simulation predicts that the clump observed in the NW tail of NGC 7252 will orbit between r ≈ 14 and 120 h^{−1} kpc with a period of 4 h^{−1} Gyr, and, if bound, it is expected to become a long-lived dwarf companion of the remnant.

(5) Is Toomre & Toomre 1972 the whole story?

The main story of Toomre & Toomre (1972) was that tails develop kinematically, and this story remains unchanged. What does need to be modified is the notion that tails are largely unbound, carrying away the excess orbital energy and angular momentum of the progenitors, allowing them to merge (§VII.b of Toomre & Toomre 1972). With the advent of modern simulation techniques and the ability to model multi-component progenitors (cf. Barnes & Hernquist 1992), it has become clear that extended dark halos can easily soak up the necessary energy and angular momentum (White 1978, Barnes 1988), and that most of the luminous matter — including the tails — will remain bound to the remnants (Barnes 1988, 1992; Hernquist 1992; Hernquist & Spergel 1992, Barnes & Hernquist 1995). The present work observationally validates this notion (§II.6.2.1, §III.4.2, Ch. V), and will hopefully add to a deeper appreciation of its implications.

First I would like to summarize my new understanding concerning the evolution of tidal material by recasting the main conclusions of Chapter IV in a somewhat different light (see also Append. VIII):

• The vast majority of the tail material remains bound to the remnant, and as such will have an associated timescale for reaching its outer radial turning point and heading towards its inner radial turning point. This is referred to as the “fall-back” time.

• Initially luminous tails are raised, but most of this material has a fall-back time that is shorter than the timescale for the potential to relax. As a result, it will fall back before the merger is complete, undergo violent relaxation, and be redistributed throughout the remnant. This material will include at least half of the progenitors disks, and therefore at least half of the original disk atomic gas supply.

• The remaining tail material (i.e. with t_{fall−back} > t_{relax}) is structured kinematically, with the most tightly bound, lowest angular momentum material at the bases of the tails and the most energetic, highest angular momentum material at the ends. This leads to a relationship between radial period and pericentric distance, both of which increase linearly with distance along the tail.

• Since this material falls back after the potential has mostly relaxed, it can wrap coherently in the potential creating fine structure features. The earliest returning material (shortest radial period) passes closest to the potential center (smallest pericentric distance) and creates shell-like features. The later returning material falls to progressively larger inner radial turning points, forming loops.

• The outermost regions of the tails evolve on very long timescales and will be distributed over a large range of radii with a low volume filling factor.

(6) Can the tail kinematics usefully constrain the encounter geometry? If so, what can we learn about both the past and future of the encounter, especially concerning the atomic gas?

Chapter IV serves as the answer to this question: given a reasonable starting point for the structure and mass ratio of the progenitors, the kinematics of the outer regions can very successfully constrain the encounter geometry. Since disk galaxies in general have an atomic-gas-to-light ratio which increases with radius (Wevers et al. 1986), and it is these outer regions which are drawn into extensive tails (§IV.5), neutral hydrogen kinematic mapping is uniquely well suited for such investigations.

In light of our new understanding concerning the development of tidal tails, it is necessary that we modify some of the existing concepts in the literature. For instance, some caution is needed when defining “tails”, since there is obviously a surface brightness involved in this definition, and the evolution of the tidal material depends critically on the surface brightness of the material you are talking about. The luminous tails which are raised shortly after orbital periapse are mostly comprised of tightly bound material that will fall back before the progenitor bodies have fully merged (§IV.4.1, esp. Fig. IV-6). The less tightly bound material will evolve on much, much longer timescales. As a result, faint tidal tails will persist well beyond the duration of the central starbursts in the ULIR systems. The fact that deep optical images of the nearby ULIR systems show them to have rather short tails (§V.2) indicates that they result from encounters which either merge very rapidly or which fail to raise significant tails (§V.3). Similarly, deep imaging of intermediate redshift samples (z_{∼}^{<}0.5) should be capable of mapping any extended tidal features (§V.3; Hibbard & Moran in preparation), whereas larger redshift samples will be sensitive to only the short-lived luminous phase of tail development (Mihos 1994). These considerations must be taken into account whenever using optically defined interacting samples to derive parent populations or probabilities.

The past evolution of the gas was covered in the answers to the first two questions: at least half of the atomic gas is converted into other phases within the remnant body during the merger. The remaining fraction is sent to large projected radii. Concerning the future evolution of this material, the simulation of NGC 7252 predicts that half of the 2×10^{9} h^{−2} M_{⊙} of tidal atomic gas will fall to within 15–45 h^{−1} kpc of the remnant over the next 3 h^{−1} Gyr. The remainder should be distributed as a very diffuse, clumpy tail with a very low column density and filling factor (§IV.4.1, Append. VIII).

The caveats attached to the beginning of this question deserve some clarification. The parameter space of mergers is orders of magnitude larger than for an individual spiral, and is very poorly understood in terms of the uniqueness of individual solutions. Therefore, at least at the present time, we must be given a reasonable model for the structure of the original galaxies. However this does not invalidate our geometric solution, since the kinematic tracers reside at large radii (r >> r_{core}) and depend primarily on the global properties of the potential (e.g. amount of dark matter), while the majority of the structure parameters (core radius, flattening, etc.) produce kinematic signatures which are most noticeable at small radii. So while it is not feasible at present to derive the details of the mass structure of the progenitors with these data, it may not be necessary or even expedient to do so.

(7) Does the combination of simulations and data give us any new information that was not already apparent in either one alone?

The answer to this question is probably no. The main benefit of marrying simulations to data is that it allows us to pay attention to details in one or the other which may not have seemed as important by themselves. For instance, the fact that tidal material remains largely bound and streams back into the remnant at late times is hardly unprecedented in the literature, but solely from the theoretical realm (Barnes 1988, 1992; Hernquist 1992; Hernquist & Spergel 1992). However, the implications of this infall was not widely appreciated. I believe this to be due in part to the lack of attention drawn to the tails in the numerical works. The comments concerning this behavior in the above references are nested within many pages of text quantifying the behavior of the remnant bodies, which contain _{∼}^{>}95% of the simulation particles, and therefore yield the best “measurements”. The tails contain only one or two thousand particles in a typical simulation, and any attempt to quantify their behavior in a statistical sense would involve large uncertainties. In Chapter V, we followed Toomre & Toomre’s (1972) lead by concentrating on the behavior of individual particles. In essence, we evolved 65,000 particles over thousands of cycles just to arrive at the orbits of 2,155 particles. Yet this seemingly wasteful exercise pays handsome dividends precisely because these 2,155 particles populate a region of the remnant that is well constrained by the observations.

Similarly, while the return of tidal material was suggested by the gas kinematics in NGC 7252 (§III.4.2), we did not realized before running the simulation that larger quantities had fallen to smaller radii in the past, and that the remaining material would fall to every more distant radii in the future (§IV.5.2). Therefore this marriage between simulations and observations proves to be happy one indeed, providing us with a very powerful analytical tool for interpreting both simulations and data in a new light. A more direct involvement between the theoretical and observational realms is one of the most promising avenues to be explored by future research.

(8) What can the observed trends and the dynamics of this sequence tell us about the connection between merging and the ULIR galaxies?

§3. Remaining Questions 183

This question was addressed in Ch. V, where I find relatively few and rather unimpressive tidal tails emanating from the luminous IR galaxies, and no ultraluminous infrared systems among the Toomre Sequence. These differences may simply reflect the short lived nature of the luminous IR phase and/or the rarity of long tails in disk-disk encounters. However they may also indicate that not all mergers of gas rich systems will experience an ultraluminous IR phase. For instance, the long luminous tails seen along the Toomre Sequence may imply somewhat wide encounters with correspondingly long merging timescales (§II.2), and it is possible that such wide encounters will not experience an ULIR phase (§V.3). Similarly, the short, luminous tails and plumes which appear in the LIR galaxies while the main bodies are already amorphous blobs may indicate small impact parameters or retrograde encounters with correspondingly short merging timescales, and such encounters may be more conducive to fueling a final super-starburst.

Alternatively, the well-defined edge-brightened appearance of the tails in the Toomre Sequence systems may indicate a more substantial low velocity dispersion disk component (i.e. gas and young stars) than for the plume-like features of the LIR systems (§II.6.2.3, §V.3). However, since the number of well studied systems is small, it hard to be conclusive, and this question should be fully investigated with larger datasets.

3. Remaining Questions

In this section I describe what I believe are the most important outstanding questions and avenues of research with respect to disk galaxy mergers.

1. The King Gap

In order to more fully constrain the evolutionary fates of these remnants, more objects in the so-called “King Gap” between merger remnants and full fledged ellipticals (§I.2.2.2) need to be studied^{1}. Based on the study in Chapter II, I expect the youngest merger remnants to have an unusual amount of molecular gas in a central star-forming disk, an intermediate-aged stellar population from the merger induced star-burst, a fair amount of optical fine structure, and relic tidal tails (§II.6.3). Since the neutral hydrogen traces the coldest component of the tails, it should be a valuable tracer of a past merger origin long after many of the stellar features have faded away, but its future appearance is hard to judge from the present work (see also Append. VIII).

With time, each of these features should fade: the molecular disk will slowly be turned into stars; the fine structure will lose its coherence after a few crossing times; the amount of material falling back from the tails will trickle off, and the outer tail will fade irregularly along its length. It is also possible that an X-ray halo is being formed at the expense of the atomic and molecular gas phases (§III.4.4). These characteristics suggest three classes of galaxies which might contain a number of aging merger remnants: early-type galaxies with unusual amounts of cold gas, galaxies with a large amount of fine structure (Schweizer & Seitzer 1992), and ellipticals with inner disk kinematics^{2}.

1 I also expect mergers to make spirals (§II.6.3.4), but until I model the NGC 520 interaction I will leave this as pure speculation.

Outer H i with interesting kinematics has been mapped in both gas-rich Es & S0s (see e.g. Appleton et al. 1990, van Driel & van Woerden 1991, and references therein) and recently in the shell galaxies (Schiminovich et al. 1994, 1995). Simulations reproducing the observed H i kinematics should be made for those systems for which a tidal origin of the gas is suspected. Similar observational/numerical investigations should be conducted on even older King Gap candidates using the dating scheme of Schweizer & Seitzer (1992). Ellipticals with inner disk kinematics or kinematically distinct sub-systems (cf. Bender 1990, Bender & Surma 1995, and references therein) may be the best candidates for the oldest remnants. It would be interesting to search for faint relic tidal appendages or other signatures of a past merger in these objects. Any candidate system should be examined for the other expected remnant characteristics, as the quantity and coherence of such signatures may be the only way to differentiate between mergers and less structurally damaging events such as multiple accretions.

2. The Fate of the Returning H i

It is still not known how the returning H i in NGC 7252 is kept from becoming distributed throughout the remnant body. At the return velocities of ∼150 km s^{−1}, this material would cross the face of the remnant in ∼10^{8} yrs, or much longer than the age of the tails. Some process must be continuously converting this gas into other phases. The H ii regions just below the dustlane as the tail crosses the face of the remnant (Ch. III Fig. 5d) suggest a shock front for this returning gas. Still, it is not clear if this front will serve mostly to compress the H i into molecular form, where it will feed the central gas disk; turn it quickly into stars; or heat it to a plasma. It is perhaps telling that the depletion rate of the infalling H i is roughly equal to current star formation rate, as inferred from the infrared and radio continuum luminosities (§III.4.4), but all of these possibilities should be investigated with detailed spectroscopic studies.

Along a similar vein, it is not clear what form the even longer period tidal material will take upon its first return. Will the gas dissipate and take up circular orbits at large radii, or sink further into the potential? Will the densities be high enough to make these features visible? Without knowing the fate of this returning material, we cannot reliably judge the future evolution of the remnants. Hydrodynamical simulations and follow-up observations of candidate remnants (i.e. King Gap objects; see above) would be instructive here.

2 I do not mean to suggest that all galaxies in these classes are the likely the result of a merger, only that these classes are good places to look for merger remnants. Clearly there are other events that would leave similar signatures. See also §4 below.

3. The Simulation/Observation Connection

The work of this thesis, as well as other H i studies (see e.g. van der Hulst 1979, Haynes et al. 1984, Rots et al. 1991, Yun et al. 1994, and references under question 3 above), demonstrate the unique ability of 21cm observations to map the distribution and kinematics of tail-like structures. In light of this ability, it would be well worth the effort to tailor the simulations to the distribution of the H i gas, which is more extended and has a shallower fall-off than the stellar light (e.g. Wevers et al. 1986, Broeils 1994). For instance, it would be reassuring to verify that the quantities of material listed under question 3 above could indeed be drawn out tidally, using progenitors with two component disks: an exponential disk with a sharp cutoff, and a flatter more extended component to represent the atomic gas. Different velocity dispersions could be assigned to each component to test whether this has any effect on the resulting tidal morphologies, as suggested in §II.6.2.3.

It would also be most desirable to have a more direct method for interfacing 3-dimensional velocity cubes obtained from spectral line mapping instruments such as the VLA, OVRO, or Fabry-Perot interferometers, to representations of the 6-dimensional simulation results. As part of this thesis, I developed software that allows a direct comparison of simulation and observational data in several 2-dimensional projections (§VI.3), but this method falls far short of fully utilizing all of the available data. These data were hard to obtain and costly to reduce, and yet much of it lies unexamined within the circuits of our workstations, representing a tremendous waste of resources. Given the three-dimensional visualization capabilities of advanced graphics workstations, is seems that such tools would be fairly straight-forward to develop, and it would be worth a major effort to do so.

4. The ULIR-Merger Connection

At the beginning of this thesis I presented the widely held paradigm that mergers of gas rich systems fuel an ultraluminous IR phase (§I.2.2.3). In the last chapter I cited references which suggest a link between such a phase and quasar activity (§V.1). So while the fraction of all galaxies which are peculiar is small (Arp 1966, Arp & Madore 1975), and the fraction which are ultraluminous in the infrared even smaller (Soifer et al. 1986), such systems may be providing important glimpses into one of the most interesting periods of a galaxy’s evolution, and may provide a local glimpse of processes which were much more common in the distant past, at the epoch of galaxy formation. It is therefore of prime importance to determine the detailed connection between these two processes — interactions/merging and elevated IR luminosity. This is an important issue to resolve if we wish to use statistics on such objects to determine the merger frequency rate or other evolutionary effects.

The outer tidal morphologies of the samples presented in Ch. V suggest a difference in either progenitor type or encounter geometry between Toomre mergers and luminous IR galaxies. In that chapter, I outlined three promising lines of research to investigate this question: (1) look for evidence for a past ULIR phase in remnants such as NGC 3921 and NGC 7252; (2) obtain an unbiased optically selected comparison sample to judge the distribution of tidal morphologies over many different encounter geometries; and (3) map the distribution and kinematics of the outer H i in the ULIR systems to directly test the geometry of the encounter, and compare these to similar results for the galaxies in the Toomre Sequence.

4. Final Comment

In this thesis I have investigated the process of merging. However, at times I have gone beyond the simple reporting of observations and indulged in the practice of fortune-telling, extrapolating the future evolution of the remnants that I study. Since I incorporate the word “fate” in the title of my thesis, I felt obliged to do this. This act involves a considerable amount of scientific bias, and for completeness I here confess to mine. I am basically driven by the following considerations:

• As documented in §I.2, we now know that a large fraction of ellipticals exhibit optical fine-structure, kinematically decoupled cores, gas which points to an external origin, and other oddities. While these signs do not necessarily point towards mergers, they do strongly suggest that the majority of galaxies do not evolve in isolation. Many of the observations can be explained most simply by gravitational interactions, with major mergers on one end of the spectrum and small accretion events on the other.

• Different types of stellar systems form specific disjoint sequences in the socalled fundamental plane of observable parameters (Djorgovski & Davis 1987). In this plane, ellipticals and the bulges of spirals form a two-parameter family in which central surface brightness decreases and core radius increases with increasing total luminosity, while pure disks, irregulars, dwarf spheriodals, and dwarf ellipticals form a separate two-parameter family in which both central surface brightness and core radius increase with increasing luminosity (cf. Kormendy & Djorgovski 1989).

• Other quantities which correlate with Hubble type that are not encompassed by the fundamental plane, such as gas content and morphology, are intricately linked to the parameter family to which the host system belongs.

So if I am to believe on the one hand that gravitation interactions are rather common, but on the other that galaxies obey certain basic relationships that constrain their structural parameters, I must believe that gravitational interactions do not introduce an unacceptable amount of scatter in the observed relationships. This implies that the relationships which differentiate galaxies of various Hubble types are set up by very general constraints (e.g., the virial equation; Faber et al. 1988), rather than on the fine details of an interaction. I find this easy to believe,

§4. Final Comment 187

since the processes involved when two galaxies merge are generic: energy dissipation, star formation, mechanical transport of angular momentum. Merging of two present-day spirals is definitely one way to get them, but undoubtably there are many others. All of these processes must have been at play whenever a galaxy formed, regardless of the formation process.

Therefore, my interest in showing that these unusual creatures will evolve to resemble their more mundane neighbors is not driven by a belief that this must be how all things are formed, but from the belief that something as common as gravitational fallback from the Hubble flow had better make normal looking things. My personal belief is that we can take this a step further and make the statement that something as common as gravitational fallback from the Hubble flow will make the whole spectrum of normal looking things: ellipticals, spirals, dwarfs. Not the only way, but a good way.

The last point implies to me that if some E’s and S0’s are anomalous in the properties of their ISM, with time they will evolve to be more representative. For instance, statistical studies of early-type galaxies suggest an external origin for any cold gas (Knapp et al. 1985, Wardle & Knapp 1986, Lees et al. 1991), which is supported by neutral hydrogen mapping in a number of gas rich systems (e.g. van Driel & van Woerden 1991, Appleton et al. 1990, and references therein). My bias given the above considerations is that these systems will either loose any outer gas (e.g. have it swept away by an IGM) or process it into stars until they resemble something more normal. Another possibility is that direct mapping of H i around otherwise normal early-type galaxies will reveal a relatively high frequency of outer H i that was missed in single dish measurements, and that cold gas in early type galaxies is much more normal than heretofore believed.

Still, we are getting pretty far afield with all this speculation. What I have said in this thesis is that galaxies can evolve from one type into another in a manner not too unlike the inverse Hubble tuning-fork diagram presented by Frano̧is Schweizer at the 1989 Heidelberg meeting (Schweizer 1990) and reproduced on the cover of this thesis: in short, I have supported the idea that two present day spirals can fall together under their mutual gravitational attraction, merge, and leave a remnant that resembles an elliptical in its photometric properties and probably also in its gas content.

This statement is different in subtle but important ways from the statement that spirals at some distant epoch can merge and evolve to resemble an old elliptical today. The kinematics of NGC 3921 and NGC 7252 imply that the interaction started around z = 0.1, while ellipticals with populations that are much older than 5 h^{−1} Gyr were formed at redshifts z > 1. We simply do not know if spirals at these high redshifts resemble those around us today. It is quite possible that the Hubble diagram evolves with redshift, implying that the appearance of future remnants may well be unlike anything we have ever seen. Thanks to a refurbished Hubble Space Telescope, we are just now getting glimpses of these distant epochs and are poised at a time in which this question can be answered.

References

van Albada, 1982, MNRAS, 201, 939. Appleton, P.N., Ghigo, F.D, van Gorkom, J.H., Schombert, J.M., Struck- Marcell, C., 1987,

Nature, 330, 140. Appleton, P.N., Pedlar, A., Wilkinson, A., 1990, ApJ, 357, 426. Arp, H.C., 1966, ApJ (Supp.), 14, 1. Arp, H., Madore, B.F., 1975, Observatory, 95, 212. Barnes, J.E., 1992, ApJ, 393, 484. Barnes, J.E., 1988, ApJ, 331, 699. Barnes, J.E., Hernquist, L., 1991, ApJ (Lett.), 370, L65. Barnes, J.E., Hernquist, L., 1992, ARAA, 30, 705. Barnes, J.E., Hernquist, L., 1995, ApJ, submitted. Bender, R., 1990, “Dynamics and Interactions of Galaxies”, ed. R. Wielen (Heidelberg: Springer),

p. 232. Bender, R., Surma, P., 1995, A&A, in press. Broeils, A.H., 1994, Ph.D. Thesis, University of Groningen. Djorgovski, S., Davis, M., 1987, ApJ, 313, 59. van Driel, W., van Woerden, H., 1991, A&A, 243, 71. Duc, P.-A, 1995, Ph. D. Thesis, Université Paris. Faber, S.M., Jackson, R., 1976, ApJ, 204, 668. Faber, S.M., Dressler, A., Davies, R.L., Burstein, D., Lynden-Bell, D., Terlevich, R., Wegner, G.,

1988, “Nearly Normal Galaxies: From the Planck Time to the Present”, ed. S.M. Faber (New York: Springer), p. 175.

Haynes, M.P., Giovanelli, R., Chincarini, G.L, 1984, ARA&A, 22, 445. Hernquist, L., 1992, ApJ, 400, 460. Hernquist, L., Spergel, D.N., 1992, ApJ, 399, L117. van der Hulst, J.M., 1979, A&A, 155, 151. Jog, C.J., Solomon, P.M., 1992, ApJ, 387, 152. Jog, C.J., Das, M., 1992, ApJ, 400, 476. Knapp, G.R., Turner, E.L., Cunniffe, P.E., 1985, AJ, 90, 454. Lake, G., Dressler, A., 1986, ApJ, 310, 605. Lees, J.F., Knapp, G.R., Rupen, M.P., Phillips, T.G., 1991, ApJ, 379, 117. Lynden-Bell, D., 1967, MNRAS, 136, 101. Mihos, J.C., 1994, ApJ (Lett.), 438, L75. Rots, A.H., Bosma, A., van der Hulst, J.M., Athanassoula, E., Crane, P.C., 1990, AJ, 100, 387. Schiminovich, D., van Gorkom, J.H., van der Hulst, J.M., Kasow, S., 1994, ApJ (Lett.), 423,

L101. Schiminovich, D., van Gorkom, J.H., van der Hulst, J.M., Malin, D.F., 1995, ApJ (Lett.), sub-

mitted. Schiminovich, D., van Gorkom, J.H., 1995, in preparation. Schneider, S.E., et al., 1989, AJ, 97, 666. Schweizer, F., 1982, ApJ, 252, 455. Schweizer, F., 1990, “Dynamics and Interactions of Galaxies”, ed. R. Wielen (Heidelberg: Springer),

p. 60. Schweizer, F., Seitzer, P., 1992, AJ, 104, 3. Simkin S.M., van Gorkom, J.H., Hibbard, J.E., Hong-Jun, S., 1987, Science, 235, 1367. Smith, B.J., 1991, ApJ, 378, 39. Smith, B.J., Higdon, J.L., 1994, AJ, in press. Soifer, B.T., Sanders, D.B., Neugebauer, G., Danielson, G.E., Lonsdale, C.J., Madore, B.F.,

Persson, S.E., 1986, ApJ (Lett.), 303, L41. Stanford, S.A., Bushouse, H.A., 1991, ApJ, 371, 92, Toomre, A., 1977, in “The Evolution of Galaxies and Stellar Populations”, eds. B.M. Tinsley

and R.B. Larson (New Haven: Yale Univ.), p. 401. Toomre, A., Toomre, J., 1972, ApJ, 178, 623. Wardle, M., Knapp, G.R., 1986, AJ, 91, 23. Wevers, B.H.M.R., van der Kruit, P.C., Allen, R.J., 1986, A&A (Supp.), 66, 505. White, S.D.M., 1978, MNRAS, 184, 185. Yun, M.S., Ho, P.T.P., Lo, K.Y., 1994, Nature, 372, 530.

Zwicky, F., 1956, Ergebnisse der Exakten Naturwissenschaften, 29, 344.